ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

19. History of Mathematics

1991 MS Classification: 1, 5, 51 Gropp, Harald, Universität Heidelberg, Germany Levi, Steinitz, Zacharias, and configurations –Three German mathematicians in the 20^{th} century and a combinatorial structure This poster discusses the lives and works of Friedrich Wilhelm Levi (1888-1966), Ernst Steinitz (1871-1928), and Max Zacharias (1873-1962). Configurations, a finite combinatorial incidence structure of points and lines, defined by Th. Reye in 1876 in a geometrical background, describe an important chapter in the history of geometry and combinatorics. Steinitz was with the TH Charlottenburg ( now TU Berlin ) between 1897 and 1910. Zacharias was a teacher in Berlin until World War II including the Nazi period, later he lived in Quedlinburg in the GDR. Levi had to leave Nazi Germany in 1935, emigrated to India, and returned to Berlin (FU) in 1952. In several aspects these three mathematicians are typical for the 20^{th} century in Germany and Berlin.